Boyong and Boyet is conducting their experiments on a frictionless surface. They observed that Boyet’s 5-kg ball travelling to the right at 2 m/s collides head-on with the stationary 7.50-kg ball of Boyong. If the collision is (a) elastic and (b) perfectly inelastic, calculate the final velocities of the balls. 


1
Expert's answer
2022-01-05T16:30:19-0500

Momentum conservation

"m_1u_1+m_2u_2=m_1v_1+m_2v_2"

Put value

"5\times2+7.50\times0=5v_1+7.50v_2"


"5v_1+7.50v_2=10"

Restitution cofficient

"e=\frac{v_2-v_1}{u_1-u_2}"

Part(a)


For elastic collision

e=1


"v_1-v_2=u_1-u_2"

"v_2-v_1=2"

Equation (1) and (2) we can solve it

"12.20v_2=20\\v_2=1.6m/sec"

Put

value in equation (2)

"v_1=0.4m\/sec"

Part(b)

Perfect inelastic

e=0

"v_2-v_1=0"

Momentum conservation

"m_1u_1+m_2u_2=m_1v_1+m_2v_2"


"5\times2+7.50\times0=5\times v_1+7.50v_2"

"5v_1+7.50v_2=10"

"v_1=v_2"

"12.50v_1=10"

"v_1=v_2=\frac{10}{12.50}=0.8m/sec"

"v_1=v_2=0.8m\/sec"